3 A car goes around a turn of radius 30 meters at a constant speed of 50 km/h. What is the acceleration of the car?

3

A car goes around a turn of radius 30 meters at a constant speed of 50 km/h. What is the acceleration of the car?

Grade:12

1 Answers

Nirmal Singh.
10 years ago
v = 50 km/ hr = 50* 5/18 m/s = 250 /18 = 125 /9 m/s
r = 30 m
acceleration a = v^2/r = *(125/9)^2 / 30 = 6.43 m/s^2
Thanks& Regards,
Nirmal Singh
Askiitians faculty